OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at C (−2, 1), D (2, 4), E (5, 0), and F (1, −3). Round your answer to the nearest hundredth when necessary. 12 units 16 units 20 units 24 units

OpenStudy (cheerleader):

I think it is 20 units

OpenStudy (cheerleader):

How i got this is i multiplied all the x's

OpenStudy (cheerleader):

So -2 * 2= 4 *5=20*1=20 so that is how i hot the answer hope its right